OpenStudy (anonymous):

~Medal+Fanning~ Henri solved the system of equations below. What mistake did he make in his work? 3x + y = 6 x − 2y = 2 y = 6 − 3x x − 2(6 − 3x) = 2 x − 12 + 6x = 2 7x − 12 = 2 + 12 +12 7 x over 7 equals 14 over 7 x = 2 3(2) + y = 6 6 + y = 6 −6 −6 y = 0 He should have substituted 6 + 3x. He combined like terms incorrectly. It should have been 5x instead of 7x. He made a mistake in distributing −2. He made no mistake.

OpenStudy (anonymous):

He made no mistake. You can substitute the value of x and y that he found look try. If x=2, y=0, 3x + y = 6 will be 3*2+0=6; x − 2y = 2 will be 2-2*0=2. So Henri made no mistake.

OpenStudy (anonymous):

There, HelloGoodmorning.

OpenStudy (anonymous):

Thank you Chineseboy :D

OpenStudy (anonymous):

You are welcome!